Jo as 37 coins(all nickles, dimes, and quarters) worth $5.50. She has 4 more quaters than nickles. How many dimes does Jo have?
Answer by stanbon(75887)

Jo as 37 coins(all nickles, dimes, and quarters) worth $5.50. She has 4 more quaters than nickles. How many dimes does Jo have?
--------------
Let number of nickels be "x"
Value of those nickels is 5x cents.
Number of quarter is "x+4"
Value of those quarters is 25(x+4)= 25x+100 cents
Number of dimes = 37-(x+x+4)=33-2x
Value of those dimes = 10(33-2x)=330-20x cents
EQUATION:
value + value + value = 550 cents
5x+25x+100+330-20x=550
10x=110
x=11 (number of nickels)
33-2x=33-22=11 (number of dimes)
Cheers,
Stan H.